# Combating Confirmation Bias: A Unified Pseudo-Labeling Framework for Entity Alignment

### v3

@@ 3.2.1 OT-based Pseudo-Labeling
Since the coupling matrix $\bm{P}^{*}\in\mathbb{R}_{+}^{|\mathcal{E}^{U}_{1}|\times|\mathcal{E}^{U}_{2}|}$ has non-negative entries, the summation $\sum_{e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}}\bm{P}^{*}_{e_{i},e_{k}}$ from Eq. (3.2.1) must be no smaller than any of its components, i.e., $\bm{P}^{*}_{e_{i},e_{k}}\leq\sum_{e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}}\bm{P}^{*}_{e_{i},e_{k}}$, $\forall e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}$. Please note that, as we focus on a transductive semi-supervised setting, the size of two unaligned entity sets, $|\mathcal{E}^{U}_{1}|$ and $|\mathcal{E}^{U}_{2}|$, are known. Together with Eq. (3.2.1), we can further derive that any component in the summation is smaller than the decision threshold, i.e., $\bm{P}^{*}_{e_{i},e_{k}}{\color[rgb]{0,0,0}\leq\sum_{e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}}\bm{P}^{*}_{e_{i},e_{k}}}<\frac{1}{2|\mathcal{E}^{U}_{1}|}$, $\forall e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}$. In other words, all other probability values in the same row of $\bm{P}^{*}_{e_{i},e_{j}}$ are smaller than the decision threshold. Thus, no conflicted entity pairs $\{(e_{i},e_{k})|\;e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}\}$ associated with $e_{i}$ are selected as pseudo-labeled alignments.
@@
Similarly, for each pseudo-labeled alignment $(e_{i},e_{j})$ with a probability value $\bm{P}^{*}_{e_{i},e_{j}}>\frac{1}{2|\mathcal{E}^{U}_{1}|}$, we can prove that no conflicted entity pairs $\{(e_{l},e_{j})|\;e_{l}\in\mathcal{E}^{U}_{1}\setminus\{e_{i}\}\}$ associated with entity $e_{j}$ are selected as pseudo-labeled alignments. Similar to Eq. (3.2.1), we can also obtain $\sum_{e_{l}\in\mathcal{E}^{U}_{1}\setminus\{e_{i}\}}\bm{P}^{*}_{e_{l},e_{j}}<\frac{1}{2|\mathcal{E}^{U}_{2}|}$ and $\bm{P}^{*}_{e_{l},e_{j}}\leq\sum_{e_{l}\in\mathcal{E}^{U}_{1}\setminus\{e_{i}\}}\bm{P}^{*}_{e_{l},e_{j}},\forall e_{l}\in\mathcal{E}^{U}_{1}\setminus\{e_{i}\}$. Together with the assumption of $|\mathcal{E}^{U}_{1}|<|\mathcal{E}^{U}_{2}|$, we can further derive that $\bm{P}^{*}_{e_{l},e_{j}}\leq\sum_{e_{l}\in\mathcal{E}^{U}_{1}\setminus\{e_{i}\}}\bm{P}^{*}_{e_{l},e_{j}}<\frac{1}{2|\mathcal{E}^{U}_{2}|}<\frac{1}{2|\mathcal{E}^{U}_{1}|}$. Therefore, all other probability values in the same column of $\bm{P}^{*}_{e_{i},e_{j}}$ are smaller than the decision threshold, i.e., $\bm{P}^{*}_{e_{l},e_{j}}<\frac{1}{2|\mathcal{E}^{U}_{1}|}$, $\forall e_{l}\in\mathcal{E}^{U}_{1}\setminus\{e_{i}\}$. Hence, no conflicted entity pairs $\{(e_{l},e_{j})|\;e_{l}\in\mathcal{E}^{U}_{1}\setminus\{e_{i}\}\}$ associated with entity $e_{j}$ are selected as pseudo-labeled alignments.
@@
In summary, we conclude that the selected pseudo-labeled alignments $\{(e_{i},e_{j})|\;\bm{P}^{*}_{e_{i},e_{j}}>\frac{1}{2\cdot\text{min}(|\mathcal{E}^{U}_{1}|,|\mathcal{E}^{U}_{2}|)},e_{i}\in\mathcal{E}^{U}_{1},e_{j}\in\mathcal{E}^{U}_{2}\}$ are guaranteed to be one-to-one alignments when $|\mathcal{E}^{U}_{1}|<|\mathcal{E}^{U}_{2}|$, and the same conclusion also holds when $|\mathcal{E}^{U}_{1}|\geq|\mathcal{E}^{U}_{2}|$. ∎
@@ 7       $\bm{P}=\bm{a}\bm{b}^{\top}\odot\bm{Q}$;
11 Select conflict-free pseudo-labeled alignments ${\color[rgb]{0,0,0}\widehat{\mathcal{S}}_{t}}=\{(e_{i},e_{j})|\;\bm{P}^{*}_{e_{i},e_{j}}>\frac{1}{2\cdot\text{min}(|\mathcal{E}^{U}_{1}|,|\mathcal{E}^{U}_{2}|)},e_{i}\in\mathcal{E}^{U}_{1},e_{j}\in\mathcal{E}^{U}_{2}\}$;
@@ 3.2.2 Parallel Pseudo-Label Ensembling
Input: Two KGs $\mathcal{G}_{1}=\{\mathcal{E}_{1},\mathcal{R}_{1},\mathcal{T}_{1}\}$, $\mathcal{G}_{2}=\{\mathcal{E}_{2},\mathcal{R}_{2},\mathcal{T}_{2}\}$, prior seed alignments $\mathcal{S}$.
@@ 4.2.1 Results on Cross-Lingual Datasets
The Results. Table 2 and Table 3 report performance comparisons on DBP15K and SRPRS, respectively. This set of results is reported with 30% prior seed alignments used for training. The asterisk (*) indicates that semantic meanings of entity names are used to construct entity features.
@@ 4.2.2 Results on Cross-Source Monolingual Datasets
The Results. Table 4 reports performance comparisons on DBP-YG-15K (OpenEA) and DBP-YG-15K (RealEA) with 30% prior seed alignments used for training. The asterisk “*” indicates that semantic meanings of entity names are used in EA modeling. Our results show that UPL-EA outperforms all five baselines across two KG pairs of DBP-YG-15K. On the OpenEA KG pair, UPL-EA achieves nearly perfect performance with all three metrics to be approximately 1, outperforming the second best baseline by more than 2% on $F_{1}$ score. On the RealEA KG pair of DBP-YG-15K, even with duplicated entities in each KG [17], UPL-EA performs competitively with an over 5% improvement on $F_{1}$ score compared to the second best baseline.
@@ 4.3.1 Effectiveness of Different Components
To assess the importance of various components of the proposed UPL-EA framework, we first conduct a thorough ablation study on five cross-lingual KG pairs from DBP15K and SRPRS. To provide deeper insights, we undertake ablation studies under two settings: the conventional setting using 30% prior seed alignments and the setting with no prior seed alignments provided. The full UPL-EA model is compared with its ablated variants, with the best performance highlighted by bold. From Table 5 and Table 6, we can see that the full UPL-EA model performs the best in all cases.
@@
Note that under the setting with no prior seed alignments, the variant without OT-based pseudo-labeling (w.o. OT. Pseudo-Labeling) has similar performance as compared to the variant completely ignoring confirmation bias (w.o. OT. & Ensemb.). In particular, on DBP15KZH_EN and DBP15KJA_EN, the former variant even performs slightly worse. This is because under the challenging case where there are no prior seed alignments, ablating OT-based pseudo-labeling might incur considerably more conflicted misalignments. As a result, it becomes ineffective to filter out erroneous pseudo-labeled alignments via ensembling.
@@ 4.3.3 Effectiveness as a General Pseudo-Labeling Framework
To further demonstrate UPL-EA’s viability as a general pseudo-labeling framework for entity alignment, we substitute the EA model in UPL-EA (described in Section 3.1) with alternative EA models, and examine if applying our UPL strategy could bring any performance improvements. We consider two alternative EA models: (1) GCN-Align [52], which adopts a two-layer GCN as an encoder to learn entity embeddings, and (2) GAT-Align, where the GCN encoder in GCN-Align is replaced with a two-layer GAT for embedding learning. Both EA models use the same loss function provided in Eq. (2). This analysis is conducted on DBP15K with 30% prior seed alignments as a case study. The entity alignment performance using the two baselines and their UPL-EA augmented counterparts is reported in Table 8.
@@
4.4 Comparison w.r.t. Different Rates of Prior Seed Alignments
Next, we further examine how the performance of UPL-EA changes with respect to different rates of prior seed alignments, decreasing from 40% to 10%. We compare UPL-EA with four representative state-of-the-art baselines (BootEA, RDGCN, RNM, and CPL-OT), and report the results on DBP15K and SRPRS in Table 9 and Table 10. The last column “$\Delta\downarrow$” in each table indicates the average performance loss when decreasing the rate from 40% to 10% for each model.
Table 9: Performance comparison (Hit@1) on DBP15K with respect to different rates of prior seed alignments. $\Delta\downarrow$” in the last column indicates the average performance loss when decreasing the rate from 40% to 10% on three datasets.
@@
Table 10: Performance comparison (Hit@1) on SRPRS with respect to different rates of prior seed alignments. “$\Delta\downarrow$” in the last column indicates the average performance loss when decreasing the rate from 40% to 10% on two datasets.
@@
As expected, UPL-EA consistently outperforms four competitors on all cross-lingual KG pairs at all prior seed alignment rates. This is due to UPL-EA’s ability to augment the training set with reliable pseudo-labeled alignments by effectively alleviating confirmation bias. As the rate of prior seed alignments decreases from 40% to 10%, the performance of BootEA significantly degrades by over 20% on average due to its limited ability to prevent the accumulation of pseudo-labeling errors. RNM outperforms RDGCN owing to its posterior embedding distance editing during pseudo-labeling; however, its lack of iterative model re-training hinders its overall performance. CPL-OT demonstrates more stable performance with varying rates of prior seed alignments because it selects pseudo-labeled alignments via the conflict-aware OT modeling and then uses them to train the EA model in turn; nevertheless, its neglect of one-to-one misalignments limits the potential of CPL-OT. UPL-EA remains consistently competitive and stable across all datasets, with an average performance loss of 1% at most when the rate of prior seed alignments decreases from 40% to 10%, even on the most challenging DBP15KZH_EN dataset.
@@ 4.5 Impact of Pre-Trained Word Embeddings
To analyze the impact of using different pre-trained word embeddings, we report the results of UPL-EA that form entity features with Glove embedding [29], which is widely used in the existing EA models. We conduct this analysis on the setting with 30% prior seed alignments. The results on DBP15K are reported in Table 11 as a case study. We can observe that UPL-EA with Glove embedding still achieves competitive results, significantly outperforming all other baselines. This confirms that the efficacy of UPL-EA is not highly dependent on embedding initialization methods used. When switching from Glove embedding to BERT pre-trained embeddings, performance gains can be observed, especially on DBP15KJA_EN. This indicates the usefulness of pre-trained word embeddings of high quality for entity alignment.
@@ 4.6 Hyper-Parameter Sensitivity Analysis
We further study the sensitivity of UPL-EA with regards to four hyper-parameters: embedding dimension $d$, number of OT-based models $M$ for pseudo-label ensembling, regularization hyper-parameter $\beta$ in Eq. (8), and margin hyper-parameter $\gamma$ in the alignment loss function Eq. (2). This set of sensitivity analysis is conducted on DBP15KZH_EN with 30% prior seed alignments as a case study. The respective results in terms of Hit@1 and Hit@10 are reported in Fig. 3.
@@ 4.7 Runtime Comparison
Lastly, we compare the overall training time of UPL-EA with three embedding-based EA models, including CPL-OT, RDGCN, and RNM, and one conventional EA model, PARIS+, across five cross-lingual datasets with 30% prior seed alignments. For a fair comparison, we use the same parameters reported in the original papers of the four baselines. Fig. 4 reports the overall training time of UPL-EA and the other baselines. Our results show that UPL-EA is considerably more efficient than the three embedding-based baselines, achieving a speedup of at least 50% across all five datasets. Although PARIS+ exhibits the shortest runtime due to its rule-based nature and lack of gradient-based optimization, UPL-EA remains highly efficient while maintaining strong EA performance.

### v4

@@ 3.2.1 OT-based Pseudo-Labeling
Since the coupling matrix $\bm{P}^{*}\in\mathbb{R}_{+}^{|\mathcal{E}^{U}_{1}|\times|\mathcal{E}^{U}_{2}|}$ has non-negative entries, the summation $\sum_{e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}}P^{*}_{e_{i},e_{k}}$ from Eq. (3.2.1) must be no smaller than any of its components, i.e., $P^{*}_{e_{i},e_{k}}\leq\sum_{e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}}P^{*}_{e_{i},e_{k}}$, $\forall e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}$. Please note that, as we focus on a transductive semi-supervised setting, the size of two unaligned entity sets, $|\mathcal{E}^{U}_{1}|$ and $|\mathcal{E}^{U}_{2}|$, are known. Together with Eq. (3.2.1), we can further derive that any component in the summation is smaller than the decision threshold, i.e., $P^{*}_{e_{i},e_{k}}{\color[rgb]{0,0,0}\leq\sum_{e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}}P^{*}_{e_{i},e_{k}}}<\frac{1}{2|\mathcal{E}^{U}_{1}|}$, $\forall e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}$. In other words, all other probability values in the same row of $P^{*}_{e_{i},e_{j}}$ are smaller than the decision threshold. Thus, no conflicted entity pairs $\{(e_{i},e_{k})|\;e_{k}\in\mathcal{E}^{U}_{2}\setminus\{e_{j}\}\}$ associated with $e_{i}$ are selected as pseudo-labeled alignments.
@@
Similarly, for each pseudo-labeled alignment $(e_{i},e_{j})$ with a probability value $P^{*}_{e_{i},e_{j}}>\frac{1}{2|\mathcal{E}^{U}_{1}|}$, we can prove that no conflicted entity pairs $\{(e_{l},e_{j})|\;e_{l}\in\mathcal{E}^{U}_{1}\setminus\{e_{i}\}\}$ associated with entity $e_{j}$ are selected as pseudo-labeled alignments. Similar to Eq. (3.2.1), we can also obtain $\sum_{e_{l}\in\mathcal{E}^{U}_{1}\setminus\{e_{i}\}}P^{*}_{e_{l},e_{j}}<\frac{1}{2|\mathcal{E}^{U}_{2}|}$ and $P^{*}_{e_{l},e_{j}}\leq\sum_{e_{l}\in\mathcal{E}^{U}_{1}\setminus\{e_{i}\}}P^{*}_{e_{l},e_{j}},\forall e_{l}\in\mathcal{E}^{U}_{1}\setminus\{e_{i}\}$. Together with the assumption of $|\mathcal{E}^{U}_{1}|<|\mathcal{E}^{U}_{2}|$, we can further derive that $P^{*}_{e_{l},e_{j}}\leq\sum_{e_{l}\in\mathcal{E}^{U}_{1}\setminus\{e_{i}\}}P^{*}_{e_{l},e_{j}}<\frac{1}{2|\mathcal{E}^{U}_{2}|}<\frac{1}{2|\mathcal{E}^{U}_{1}|}$. Therefore, all other probability values in the same column of $P^{*}_{e_{i},e_{j}}$ are smaller than the decision threshold, i.e., $P^{*}_{e_{l},e_{j}}<\frac{1}{2|\mathcal{E}^{U}_{1}|}$, $\forall e_{l}\in\mathcal{E}^{U}_{1}\setminus\{e_{i}\}$. Hence, no conflicted entity pairs $\{(e_{l},e_{j})|\;e_{l}\in\mathcal{E}^{U}_{1}\setminus\{e_{i}\}\}$ associated with entity $e_{j}$ are selected as pseudo-labeled alignments.
@@
In summary, we conclude that the selected pseudo-labeled alignments $\{(e_{i},e_{j})|\;P^{*}_{e_{i},e_{j}}>\frac{1}{2\cdot\text{min}(|\mathcal{E}^{U}_{1}|,|\mathcal{E}^{U}_{2}|)},e_{i}\in\mathcal{E}^{U}_{1},e_{j}\in\mathcal{E}^{U}_{2}\}$ are guaranteed to be one-to-one alignments when $|\mathcal{E}^{U}_{1}|<|\mathcal{E}^{U}_{2}|$, and the same conclusion also holds when $|\mathcal{E}^{U}_{1}|\geq|\mathcal{E}^{U}_{2}|$. ∎
@@ 7       $\bm{P}=\bm{a}\bm{b}^{\top}\odot\bm{Q}$;
11 Select conflict-free pseudo-labeled alignments ${\color[rgb]{0,0,0}\widehat{\mathcal{S}}_{t}}=\{(e_{i},e_{j})|\;P^{*}_{e_{i},e_{j}}>\frac{1}{2\cdot\text{min}(|\mathcal{E}^{U}_{1}|,|\mathcal{E}^{U}_{2}|)},e_{i}\in\mathcal{E}^{U}_{1},e_{j}\in\mathcal{E}^{U}_{2}\}$;
@@ 3.2.2 Parallel Pseudo-Label Ensembling
Input: Two KGs $\mathcal{G}_{1}=\{\mathcal{E}_{1},\mathcal{R}_{1},\mathcal{T}_{1}\}$, $\mathcal{G}_{2}=\{\mathcal{E}_{2},\mathcal{R}_{2},\mathcal{T}_{2}\}$, seed alignments $\mathcal{S}$.
@@ 4.2.1 Results on Cross-Lingual Datasets
The Results. Table 2 and Table 3 report performance comparisons on DBP15K and SRPRS, respectively. This set of results is reported with 30% seed alignments used for training. The asterisk (*) indicates that semantic meanings of entity names are used to construct entity features.
@@ 4.2.2 Results on Cross-Source Monolingual Datasets
The Results. Table 4 reports performance comparisons on DBP-YG-15K (OpenEA) and DBP-YG-15K (RealEA) with 30% seed alignments used for training. The asterisk “*” indicates that semantic meanings of entity names are used in EA modeling. Our results show that UPL-EA outperforms all five baselines across two KG pairs of DBP-YG-15K. On the OpenEA KG pair, UPL-EA achieves nearly perfect performance with all three metrics to be approximately 1, outperforming the second best baseline by more than 2% on $F_{1}$ score. On the RealEA KG pair of DBP-YG-15K, even with duplicated entities in each KG [17], UPL-EA performs competitively with an over 5% improvement on $F_{1}$ score compared to the second best baseline.
@@ 4.3.1 Effectiveness of Different Components
To assess the importance of various components of the proposed UPL-EA framework, we first conduct a thorough ablation study on five cross-lingual KG pairs from DBP15K and SRPRS. To provide deeper insights, we undertake ablation studies under two settings: the conventional setting using 30% seed alignments and the setting with no seed alignments provided. The full UPL-EA model is compared with its ablated variants, with the best performance highlighted by bold. From Table 5 and Table 6, we can see that the full UPL-EA model performs the best in all cases.
@@
Note that under the setting with no seed alignments, the variant without OT-based pseudo-labeling (w.o. OT. Pseudo-Labeling) has similar performance as compared to the variant completely ignoring confirmation bias (w.o. OT. & Ensemb.). In particular, on DBP15KZH_EN and DBP15KJA_EN, the former variant even performs slightly worse. This is because under the challenging case where there are no seed alignments, ablating OT-based pseudo-labeling might incur considerably more conflicted misalignments. As a result, it becomes ineffective to filter out erroneous pseudo-labeled alignments via ensembling.
@@ 4.3.3 Effectiveness as a General Pseudo-Labeling Framework
To further demonstrate UPL-EA’s viability as a general pseudo-labeling framework for entity alignment, we substitute the EA model in UPL-EA (described in Section 3.1) with alternative EA models, and examine if applying our UPL strategy could bring any performance improvements. We consider two alternative EA models: (1) GCN-Align [52], which adopts a two-layer GCN as an encoder to learn entity embeddings, and (2) GAT-Align, where the GCN encoder in GCN-Align is replaced with a two-layer GAT for embedding learning. Both EA models use the same loss function provided in Eq. (2). This analysis is conducted on DBP15K with 30% seed alignments as a case study. The entity alignment performance using the two baselines and their UPL-EA augmented counterparts is reported in Table 8.
@@
4.4 Comparison w.r.t. Different Rates of Seed Alignments
Next, we further examine how the performance of UPL-EA changes with respect to different rates of seed alignments, decreasing from 40% to 10%. We compare UPL-EA with four representative state-of-the-art baselines (BootEA, RDGCN, RNM, and CPL-OT), and report the results on DBP15K and SRPRS in Table 9 and Table 10. The last column “$\Delta\downarrow$” in each table indicates the average performance loss when decreasing the rate from 40% to 10% for each model.
Table 9: Performance comparison (Hit@1) on DBP15K with respect to different rates of seed alignments. $\Delta\downarrow$” in the last column indicates the average performance loss when decreasing the rate from 40% to 10% on three datasets.
@@
Table 10: Performance comparison (Hit@1) on SRPRS with respect to different rates of seed alignments. “$\Delta\downarrow$” in the last column indicates the average performance loss when decreasing the rate from 40% to 10% on two datasets.
@@
As expected, UPL-EA consistently outperforms four competitors on all cross-lingual KG pairs at all seed alignment rates. This is due to UPL-EA’s ability to augment the training set with reliable pseudo-labeled alignments by effectively alleviating confirmation bias. As the rate of seed alignments decreases from 40% to 10%, the performance of BootEA significantly degrades by over 20% on average due to its limited ability to prevent the accumulation of pseudo-labeling errors. RNM outperforms RDGCN owing to its posterior embedding distance editing during pseudo-labeling; however, its lack of iterative model re-training hinders its overall performance. CPL-OT demonstrates more stable performance with varying rates of seed alignments because it selects pseudo-labeled alignments via the conflict-aware OT modeling and then uses them to train the EA model in turn; nevertheless, its neglect of one-to-one misalignments limits the potential of CPL-OT. UPL-EA remains consistently competitive and stable across all datasets, with an average performance loss of 1% at most when the rate of seed alignments decreases from 40% to 10%, even on the most challenging DBP15KZH_EN dataset.
@@ 4.5 Impact of Pre-Trained Word Embeddings
To analyze the impact of using different pre-trained word embeddings, we report the results of UPL-EA that form entity features with Glove embedding [29], which is widely used in the existing EA models. We conduct this analysis on the setting with 30% seed alignments. The results on DBP15K are reported in Table 11 as a case study. We can observe that UPL-EA with Glove embedding still achieves competitive results, significantly outperforming all other baselines. This confirms that the efficacy of UPL-EA is not highly dependent on embedding initialization methods used. When switching from Glove embedding to BERT pre-trained embeddings, performance gains can be observed, especially on DBP15KJA_EN. This indicates the usefulness of pre-trained word embeddings of high quality for entity alignment.
@@ 4.6 Hyper-Parameter Sensitivity Analysis
We further study the sensitivity of UPL-EA with regards to four hyper-parameters: embedding dimension $d$, number of OT-based models $M$ for pseudo-label ensembling, regularization hyper-parameter $\beta$ in Eq. (8), and margin hyper-parameter $\gamma$ in the alignment loss function Eq. (2). This set of sensitivity analysis is conducted on DBP15KZH_EN with 30% seed alignments as a case study. The respective results in terms of Hit@1 and Hit@10 are reported in Fig. 3.
@@ 4.7 Runtime Comparison
Lastly, we compare the overall training time of UPL-EA with three embedding-based EA models, including CPL-OT, RDGCN, and RNM, and one conventional EA model, PARIS+, across five cross-lingual datasets with 30% seed alignments. For a fair comparison, we use the same parameters reported in the original papers of the four baselines. Fig. 4 reports the overall training time of UPL-EA and the other baselines. Our results show that UPL-EA is considerably more efficient than the three embedding-based baselines, achieving a speedup of at least 50% across all five datasets. Although PARIS+ exhibits the shortest runtime due to its rule-based nature and lack of gradient-based optimization, UPL-EA remains highly efficient while maintaining strong EA performance.